WORK AND POWER PRACTICE PROBLEMS

  1. John pushes horizontally with an 40 N force on a 50 kg mass 20 m across a floor Calculate the amount of work lee did If this task takes 1.0 minutes, how much power was generated?

  2. The third floor of a house is 9.0 m above street level How much week is needed to move a 175 kg refrigerator from the street to the third floor? If this takes 300 0 s. how much power was generated?

  3. Rebecca pulls a sled across level snow with a force of 125 N along a rope that 50
Answer #1

Number 1)

W = Fd = (40)(20) = 800 J

P = W/t = 800/60 = 13.3 W

Number 2)

W = (175)(9.8)(9) = 15435 J

P = 15435/300 = 51.45 W

Number 3)

W = Fdcos(angle)

W = (125)(55)(cos 50) = 4419 J

Number 4)

W = Fdcos(angle)

1.1 X 104 = (125)(155)(cos angle)

angle = 55.4o

5) This is 0 Joules. The force was upward and distance horizontal, no work done

6) P = Fv = (7000)(25) = 1.75 X 105 W

7) P = Fv

190 = F(.06)

F = 3167 N